# A remark on inverse limits of effective subshifts

### v1

@@ 1 Introduction
Effective actions can be formally defined using computable analysis. We say that an action of a finitely generated group $G\curvearrowright X$ is an effective dynamical system (EDS), if it is topologically conjugate to an action of $G$ on a recursively compact subset of a computable metric space, in such a way that every generator of $G$ acts through computable maps. This class of actions is intimately related to other natural classes of group actions, such as subshifts of finite type, see [8, 3]. The understanding of this relationship has led to interesting results, such as the classification of the numbers that arise as topological entropies of subshifts of finite type [9] in $\mathbb{Z}^{2}$ and the existence of chaotic behavior at zero-temperature for locally constant potentials [6].
While the class of EDS satisfies several interesting dynamical properties, in general it is not closed under topological factor maps (see the examples in [3, Propositions 8.1 and 8.2]). However, the class of topological factors of an EDS does satisfy a weak computability constraint: it must be topologically conjugate to the inverse limit of a (non-necessarily uniform) sequence of effective subshifts [3, Corollary 8.8]. This led the authors of [3] to define the class of weakly effective dynamical systems (wEDS) as that of the groups actions $G\curvearrowright X$ which are topologically conjugate to the inverse limit of a sequence of effective $G$-subshifts. The class of wEDS is stable under topological factor maps, and every EDS is a wEDS.
The main motivation behind the name wEDS in [3] was an initial belief that they might characterize the class of topological factors of EDS ([3, Question 8.12]). Our main result is that in fact there exists wEDS which are not the topological factor of any EDS in a large class of groups.
@@
The domino problem of a finitely generated group $G$ is undecidable if there is no algorithm which, given as input a finite set of forbidden patterns, can decide whether the subshift of finite type induced by those patterns is nonempty. This notion is based on the classical result by Berger [4] that states, in modern terms, that the domino problem is undecidable for $G=\mathbb{Z}^{2}$. We refer the reader to [1] for a survey on the domino problem for groups, and note that it has been conjectured that the only groups for which the domino problem is decidable are the virtually free groups [2].
@@
The first a ingredient is a characterization, initially proved for $\mathbb{Z}^{d}$-actions by Hochman [7], of the groups with decidable word problem which admit a universal EDS.
@@
We remark that B is just the natural extension of Theorems 1.1 and 1.2 of [7] to finitely generated groups with decidable word problem, and that this generalization does not present any additional difficulties. We provide this proof in Section 3 for the sake of completeness.
@@ Acknowledgments
The authors are grateful to N. Carrasco-Vargas for suggesting helpful improvements on our initial draft. S. Barbieri was supported by the ANID project FONDECYT regular 1240085, AMSUD240026 and ECOS230003. L. Poirier acknowledges that this research was partly done in the University of Santiago de Chile during an internship and is grateful to ENS de Lyon for their funding and support.
@@ 2.1 Subshifts and morphisms
Fix a finite set $A$. Given a finite $F\subset G$, a map $p\colon F\to A$ is called a pattern with support $F$ and its associated cylinder set is given by $[p]=\{x\in A^{G}:x|_{F}=p\}$. A subshift can be equivalently defined as a subset of $A^{G}$ which avoids a collection of patterns. That is, for which there exists a collection $\mathcal{F}$ of patterns such that
@@
We shall need the following characterization of morphisms between subshifts, for a proof see [5, Theorem 1.8.1].
@@ 2.2 The domino problem
We note that for every recursively presented and finitely generated group there exists an algorithm which on input a finite set of patterns $\mathcal{F}$ halts if and only if the subshift $X_{\mathcal{F}}$ is empty (see [1, Proposition 9.3.29]). Hence we may alternatively say that $G$ has decidable domino problem, if there is an algorithm which on input a finite set of patterns $\mathcal{F}$ halts if and only if the subshift $X_{\mathcal{F}}$ is nonempty.
We also note that to our current knowledge, the only infinite groups known to have decidable domino problem are the virtually free groups. It has been conjectured that they are the only ones [2].
@@ 2.3 Computability in zero-dimensional spaces
Next we shall succinctly introduce computability notions for subsets of the Cantor space. For a friendlier presentation we refer the reader to [3].
@@
Let $A,B$ be finite sets. Given $X\subset A^{\mathbb{N}}$, we say a map $f\colon X\to B^{\mathbb{N}}$ is computable if there exists a partial computable map $g\colon A^{*}\to B^{*}$ such that for any $x\in X$, $g(x|_{\{0,\dots,n\}})$ is defined for every $n$, the cylinders $[g(x|_{\{0,\dots,n\}})]$ form a nested sequence and
@@
We note that there is a more general definition of EDS on which $G$ is allowed to act in more general metric spaces (see [3, Definition 3.25]), however, the definitions are equivalent in the zero-dimensional case [3, Proposition 4.10], and furthermore, for recursively presented groups (in particular, for groups with decidable word problem), every EDS is the topological factor of a zero-dimensional EDS [3, Theorem A]. Therefore, for the purpose of this note, the zero-dimensional definition given above is enough.
@@
We note that in the case where $G$ is a finitely generated group with decidable word problem, we may identify patterns $p\colon F\to A$ with words in $A^{*}$ and thus speak about recursively enumerable sets of patterns. The following characterization of effective subshifts is well-known, a proof can be found in [3, Corollary 7.7].
@@
Notice that the left shift action $G\curvearrowright A^{G\times\mathbb{N}}$ is computable. The following characterization of zero-dimensional EDS is quite useful as it encodes the computability of the action in the space. We give a brief proof sketch.
@@ 2.4 Weak effective dynamical systems and universality
An action of a group $G$ on a zero-dimensional space is called a weak effective dynamical system (wEDS) if it is topologically conjugate to an inverse limit of effective subshifts.
We note that, as a consequence of Proposition 2.9 every zero-dimensional EDS can be represented as an inverse limit of effective subshifts, however, it has the added property that the sequence of effective subshifts $(X_{n})_{n\in\mathbb{N}}$ is uniform, that is, there is a single algorithm which on input $n$ produces a list of forbidden patterns for $X_{n}$. In our definition of wEDS we do not ask for the sequence to be uniform.
@@ 3 Existence of universal EDS
Let $(L_{n})_{n\in\mathbb{N}}$ be a recursive enumeration of all recursively enumerable subsets of patterns of the form $\{0,1\}^{F}$ for some finite $F\subset G\times\mathbb{N}$. In this way, every zero-dimensional EDS for some $n\in\mathbb{N}$ is topologically conjugate to the shift action on
@@
Consider a recursive enumeration $p_{0},p_{1},\dots$ of all patterns of the form $\{0,1\}^{F}$ for some finite $F\subset G\times\mathbb{N}$. For each $n\in\mathbb{N}$, we construct a sublanguage $L^{\prime}_{n}\subset L_{n}$ as follows:
@@
Assume without loss of generality that $\mathfrak{U}\subset\{0,1\}^{G\times\mathbb{N}}$ and that $G\curvearrowright\mathfrak{U}$ is the left shift map. Let $(u_{i})_{i\in\mathbb{N}}$ be a recursively enumerable sequence of maps from a finite subset of $G\times\mathbb{N}$ to $\{0,1\}$ such that
@@
As $(C^{t}_{i})_{i=1}^{n}$ is a partition of $\{0,1\}^{\mathbb{N}}$, the map $f$ is well defined and continuous. Furthermore, if we equip $A^{G}$ with the shift action, we have that $f$ is $G$-equivariant, namely, for every $g\in G$ we have that $f(gx)=g(f(x))$. Moreover, as $\mathfrak{U}$ is nonempty, we have that $f(\mathfrak{U})$ is a nonempty $G$-subshift. Finally, as $\mathfrak{U}\subset\mathfrak{U}_{m}$ and $P_{m,t}\cap\mathcal{F}=\varnothing$, we obtain that for each $x\in\mathfrak{U}$, we have $f(x)|_{W}\notin\mathcal{F}$, and thus $f(\mathfrak{U})\subset X_{\mathcal{F}}$, which is thus nonempty.
@@
$\mathfrak{U}_{m}\cap\bigcap_{g\in W}g^{-1}[w^{t}_{\varphi(g)}]=\varnothing\mbox{ if and only if }\mathfrak{U}\cap\bigcap_{g\in W}g^{-1}[w^{t}_{\varphi(g)}]=\varnothing.$
Hence if $\varphi\in P_{m,t}$, we have that for $x\in\mathfrak{U}\cap\bigcap_{g\in W}g^{-1}C^{t}_{\varphi(g)}=\bigcap_{g\in W}g^{-1}U_{i}$ we get $f(x)|_{W}=\varphi$. Furthermore, as $f(x)\in X_{\mathcal{F}}$, we get that $f(x)|_{W}\notin\mathcal{F}$ and thus we deduce that $P_{m,t}\cap\mathcal{F}=\varnothing$.
@@ 4 Existence of universal wEDS
Consider the collection of all tuples of the form $(h,s,X,Y,f)$ where $h$ and $s$ are non-negative integers, $X$ and $Y$ are nonempty effective $G$-subshifts whose alphabets are finite subsets of $\mathbb{N}$ and $f$ is a topological factor map between $X$ and $Y$.
There are at most countably many effective $G$-subshifts (with their alphabets identified with finite subset of $\mathbb{N}$) for any given group $G$. Moreover, by the Curtis-Hedlund-Lyndon theorem, there can be at most countably many topological morphisms between $X$ and $Y$ and thus at most countably many topological factor maps. Therefore the space of all such tuples is countable.
@@
Let us argue that $\mathfrak{X}$ is a wEDS. Consider the projection map $\pi_{k}\colon\mathfrak{X}\to\prod_{n\in A_{k}}X_{n}$ given by $\pi_{k}((x_{n})_{n\in\mathbb{N}})=(x_{n})_{n\in A_{k}}$ and let $\mathfrak{X}_{k}=\pi_{k}(\mathfrak{X})$. It is clear that $\mathfrak{X}_{k}$ is a subshift as it is a closed $G$-invariant subset of $\prod_{n\in A_{k}}X_{k}$. Furthermore, if we define for $i\in\mathbb{N}$ the natural projection map $g_{i}\colon\mathfrak{X}_{i+1}\to\mathfrak{X}_{i}$, then $\mathfrak{X}$ is the inverse limit of the sequence $(\mathfrak{X}_{k})_{k\in\mathbb{N}}$ endowed with the collection of factor maps $g_{i}$. Therefore, in order to show that $\mathfrak{X}$ is a wEDS, we only need to argue that for every $k\geq 0$, the subshift $\mathfrak{X}_{k}$ is effective. Indeed, let $k\geq 0$ and note first that $\prod_{n\in A_{k}}X_{k}$ is a nonempty product of effective subshifts and thus a nonempty effective subshift itself. Furthermore, by our definition of $A_{k}$, it follows that $\mathfrak{X}_{k}$ is precisely the set of elements of $\prod_{n\in A_{k}}X_{k}$ which satisfy conditions $\mathtt{C}_{n}$ for each $n\in A_{k}$. Noting that each of these conditions can be implemented with finitely many forbidden patterns (as a consequence again of the Curtis-Hedlund-Lyndon theorem), we conclude that $\mathfrak{X}_{k}$ is effective.
Finally, let us show that $G\curvearrowright\mathfrak{X}$ is universal. Let $G\curvearrowright X$ be a wEDS, then it is topologically conjugate to the inverse limit of a sequence of effective $G$-subshifts. Let $(Y_{i})_{i\in\mathbb{N}}$ be said sequence of effective subshifts and for $i\in\mathbb{N}$ let $t_{i}\colon Y_{i+1}\to Y_{i}$ be the associated factor maps.
@@
    If $i<H(n)$, we choose $x_{D_{i}(n)}\in X_{D_{i}(n)}$ such that $f_{D_{i}(n)}(x_{D_{i}(n)})=x_{D_{i+1}(n)}$ and repeat the process with $i-1$ until all the values $x_{D_{i}(n)}$ are defined.
@@
Let us now consider the situation of a group $G$ which does not have decidable word problem. In this general setting, we may no longer characterize effective subshifts through forbidden patterns on $G$, but we may identify them in such a way lifting the action of $G$ to a free group [3, Proposition 7.1]. More precisely, if we let $S\subset G$ be a finite set of generators and consider the free group $F(S)$ generated by $S$ and the canonical epimorphism $\phi\colon F(S)\to G$, then a subshift $X\subset A^{G}$ is effective if and only if its pullback $\widehat{X}\subset A^{F(S)}$ is effective, where
@@
Thus in the proof of Theorem 4.1 we may consider instead of $G$ the group $F(S)$ and instead of effective $G$-subshifts $X_{n}$ and $Y_{n}$ we replace them by their pullbacks $\widehat{X}_{n}$ and $\widehat{Y}_{n}$. The rest of the proof is exactly the same and at the end we note that the kernel of the action $F(S)\curvearrowright\mathfrak{X}$ contains $\operatorname{ker}(\phi)$ and thus the resulting induced action of $G$ is the universal $G$-wEDS. We thus obtain C.

### v2

@@ 1 Introduction
Effective actions can be formally defined using computable analysis. We say that an action of a finitely generated group $G\curvearrowright X$ is an effective dynamical system (EDS), if it is topologically conjugate to an action of $G$ on a recursively compact subset of a computable metric space, in such a way that every generator of $G$ acts through computable maps. This class of actions is intimately related to other natural classes of group actions, such as subshifts of finite type, see [Hochman2009b_simulation, BarCarRoj2025]. The understanding of this relationship has led to interesting results, such as the classification of the numbers that arise as topological entropies of subshifts of finite type [HochmanMeyerovitch2010] in $\mathbb{Z}^{2}$ and the existence of chaotic behavior at zero-temperature for locally constant potentials [Chazottes2010].
While the class of EDS satisfies several interesting dynamical properties, in general it is not closed under topological factor maps (see the examples in [BarCarRoj2025, Propositions 8.1 and 8.2]). However, the class of topological factors of an EDS does satisfy a weak computability constraint: it must be topologically conjugate to the inverse limit of a (non-necessarily uniform) sequence of effective subshifts [BarCarRoj2025, Corollary 8.8]. This led the authors of [BarCarRoj2025] to define the class of weakly effective dynamical systems (wEDS) as that of the groups actions $G\curvearrowright X$ which are topologically conjugate to the inverse limit of a sequence of EDS, or equivalently (see Proposition 2.11), to an inverse limit of effective subshifts. The class of wEDS is stable under topological factor maps, and every EDS is a wEDS.
The main motivation behind the name wEDS in [BarCarRoj2025] was an initial belief that they might characterize the class of topological factors of EDS ([BarCarRoj2025, Question 8.12]). Our main result is that in fact there exists wEDS which are not the topological factor of any EDS in a large class of groups.
@@
The domino problem of a finitely generated group $G$ is undecidable if there is no algorithm which, given as input a finite set of forbidden patterns, can decide whether the subshift of finite type induced by those patterns is nonempty. This notion is based on the classical result by Berger [berger_undecidability_1966] that states, in modern terms, that the domino problem is undecidable for $G=\mathbb{Z}^{2}$. We refer the reader to [ABJ2018] for a survey on the domino problem for groups, and note that it has been conjectured that the only groups for which the domino problem is decidable are the virtually free groups [ballier_domino_2018].
@@
The first a ingredient is a characterization, initially proved for $\mathbb{Z}^{d}$-actions by Hochman [Hochman2009], of the groups with decidable word problem which admit a universal EDS.
@@
We remark that B is just the natural extension of Theorems 1.1 and 1.2 of [Hochman2009] to finitely generated groups with decidable word problem, and that this generalization does not present any additional difficulties. We provide this proof in Section 3 for the sake of completeness.
@@ Acknowledgments
The authors are grateful to N. Carrasco-Vargas for suggesting helpful improvements on our initial draft. We also thank an anonymous referee for their suggestions which greatly improved the exposition of this note. S. Barbieri was supported by the ANID project FONDECYT regular 1240085, AMSUD240026 and ECOS230003. L. Poirier acknowledges that this research was partly done in the University of Santiago de Chile during an internship and is grateful to ENS de Lyon for their funding and support.
@@ 2.1 Subshifts and morphisms
Fix a finite set $A$. Given a finite $F\subset G$, a map $p\colon F\to A$ is called a pattern with support $F$ and its associated cylinder set is given by $p]=\{x\in A^{G}:x|_{F}=p\}$. A subshift can be equivalently defined as a subset of $A^{G}$ which avoids a collection of patterns. That is, for which there exists a collection $\mathcal{F}$ of patterns such that
@@
We shall need the following characterization of morphisms between subshifts, for a proof see [ceccherini2010cellular, Theorem 1.8.1].
@@ 2.2 The domino problem
We note that for every recursively presented and finitely generated group there exists an algorithm which on input a finite set of patterns $\mathcal{F}$ halts if and only if the subshift $X_{\mathcal{F}}$ is empty (see [ABJ2018, Proposition 9.3.29]). Hence we may alternatively say that $G$ has decidable domino problem, if there is an algorithm which on input a finite set of patterns $\mathcal{F}$ halts if and only if the subshift $X_{\mathcal{F}}$ is nonempty.
We also note that to our current knowledge, the only finitely generated groups known to have decidable domino problem are the virtually free groups. It has been conjectured that they are the only ones [ballier_domino_2018].
@@ 2.3 Computability in zero-dimensional spaces
Next we shall succinctly introduce computability notions for subsets of the Cantor space. For a friendlier presentation we refer the reader to [BarCarRoj2025].
@@
Let $A,B$ be finite sets. Given $X\subset A^{\mathbb{N}}$, we say a map $f\colon X\to B^{\mathbb{N}}$ is computable if there exists a partial computable map $g\colon A^{*}\to B^{*}$ such that for any $x\in X$, $g(x|_{\{0,\dots,n\}})$ is defined for every $n$, the cylinders $g(x|_{\{0,\dots,n\}})]$ form a nested sequence and
@@
We note that there is a more general definition of EDS on which $G$ is allowed to act in more general metric spaces (see [BarCarRoj2025, Definition 3.25]), however, the definitions are equivalent in the zero-dimensional case [BarCarRoj2025, Proposition 4.10], and furthermore, for recursively presented groups (in particular, for groups with decidable word problem), every EDS is the topological factor of a zero-dimensional EDS [BarCarRoj2025, Theorem A]. Therefore, for the purpose of this note, the zero-dimensional definition given above is enough.
@@
We recall that in the case where $G$ is a finitely generated group with decidable word problem we have a natural bijection $\nu\colon\mathbb{N}\to G$ that makes the standard group operations computable. Using this, we may encode a pattern $p\colon F\to A$ where $F\subset G$ is finite as a word $w\in(A\cup\{\times\})^{*}$, where the word $w=w_{0}\dots w_{n-1}$ codifies the pattern $p_{w}$ with support $F_{w}=\{\nu(i):i\in\{0,\dots,n-1\}\mbox{ and }w_{i}\neq\times\}$ and is defined by $p_{w}(\nu(i))=w_{i}$. It follows that we may unambiguously speak about recursively enumerable sets of patterns in such groups. The following characterization of effective subshifts is well-known, a proof can be found in [BarCarRoj2025, Corollary 7.7].
@@
Notice that the left shift action $G\curvearrowright A^{G\times\mathbb{N}}$ is computable. The following characterization of zero-dimensional EDS is quite useful as it encodes the computability of an arbitrary action in the topology of the space. We give a brief proof sketch.
@@ 2.4 Weak effective dynamical systems and universality
An action of a group $G$ on a zero-dimensional space is called a weak effective dynamical system (wEDS) if it is topologically conjugate to an inverse limit of EDS.
We remark that in [BarCarRoj2025] a wEDS is defined as the inverse limit of effective subshifts rather than of arbitrary EDS. These two notions are equivalent.
Proposition 2.11.
A $G$-action is a wEDS if and only if it topologically conjugate to an inverse limit of effective subshifts.
Proof.
Effective subshifts are EDS, so naturally every inverse limit of subshifts is a wEDS. For the converse, we employ a diagonal argument. Let $(G\curvearrowright X_{n})_{n\in\mathbb{N}}$ be a sequence of EDS. By Proposition 2.9, we can without loss of generality assume $G\curvearrowright X_{n}$ is the shift action on an effectively closed and shift invariant $X_{n}\subset\{0,1\}^{G\times\mathbb{N}}$. Let $\pi_{n}\colon X_{n+1}\to X_{n}$ be factor maps and let $X$ be the inverse limit associated to these sequences.
Consider $Z\subset\{0,1\}^{G\times\mathbb{N}^{2}}$ as the set of maps $z\colon G\times\mathbb{N}^{2}\to\{0,1\}$ such that if we let $x_{n}\in\{0,1\}^{G\times\mathbb{N}}$ be given by $x_{n}(g,i)=z(g,i,n)$, then for each $n\in\mathbb{N}$ we have $x_{n}\in X_{n}$ and $\pi_{n}(x_{n+1})=x_{n}$. Endow $Z$ with the natural shift $G$-action. It is clear from the definition that $G\curvearrowright Z$ is topologically conjugate to the endowed pointwise $G$-action on $X$.
For each $t\in\mathbb{N}$, define $Y_{t}$ as the set
$Y_{t}$ $=\{y\colon G\to\{0,1\}^{\{0,\dots,t\}^{2}}:\mbox{ there is }z\in Z:\bigl(y(g)\bigr)(i,n)=z(g,i,n)\}$
$=\{y\colon G\to\{0,1\}^{\{0,\dots,t\}^{2}}:\mbox{ there is }x\in X_{t+1}:\bigl(y(g)\bigr)(i,n)=\bigl(\pi_{n}\circ\dots\circ\pi_{t}(x)\bigr)(g,i)\}$
From the second equality and the fact that $X_{t+1}$ is an EDS it follows that $Y_{t}$ is effectively closed. As $Z$ is shift invariant, it follows that so is $Y_{t}$ and thus the shift action $G\curvearrowright Y_{t}$ is an effective subshift. Finally, if we endow the sequence $(G\curvearrowright Y_{t})$ with the natural factor maps that restrict the maps in each coordinate, we have that the pointwise shift $G$-action on the inverse limit $\lim_{\leftarrow}Y_{t}$ is topologically conjugate to $G\curvearrowright Z$.∎
We note that, as a consequence of Proposition 2.9 every zero-dimensional EDS can be represented as an inverse limit of effective subshifts, however, it has the added property that the sequence of effective subshifts $(X_{n})_{n\in\mathbb{N}}$ is uniform, that is, there is a single algorithm which on input $n$ produces a list of forbidden patterns for $X_{n}$ and its associated factor map. In our definition of wEDS we do not ask for the sequences to be uniform.
@@ 3 Existence of universal EDS
Let $(L_{n})_{n\in\mathbb{N}}$ be a recursive enumeration of all recursively enumerable subsets of patterns of the form $\{0,1\}^{F}$, where $F$ ranges over all finite subsets of $G\times\mathbb{N}$. In this way, every zero-dimensional EDS for some $n\in\mathbb{N}$ is topologically conjugate to the shift action on
@@
Consider a recursive enumeration $p_{0},p_{1},\dots$ of all patterns of the form $\{0,1\}^{F}$, where $F$ ranges over all finite subsets of $G\times\mathbb{N}$. For each $n\in\mathbb{N}$, we construct a sublanguage $L^{\prime}_{n}\subset L_{n}$ as follows:
@@
Assume without loss of generality that $\mathfrak{U}\subset\{0,1\}^{G\times\mathbb{N}}$ and that $G\curvearrowright\mathfrak{U}$ is the left shift map. Let $(u_{i})_{i\in\mathbb{N}}$ be a recursively enumerable sequence of maps from finite subsets of $G\times\mathbb{N}$ to $\{0,1\}$ such that
@@
As $(C^{t}_{i})_{i=1}^{n}$ is a partition of $\{0,1\}^{\mathbb{N}}$, the map $f$ is well defined and continuous. Furthermore, if we equip $A^{G}$ with the shift action, we have that $f$ is $G$-equivariant, namely, for every $g\in G$ we have that $f(gx)=g(f(x))$. Moreover, as $\mathfrak{U}$ is nonempty, we have that $f(\mathfrak{U})$ is a nonempty subshift. Finally, as $\mathfrak{U}\subset\mathfrak{U}_{m}$ and $P_{m,t}\cap\mathcal{F}=\varnothing$, we obtain that for each $x\in\mathfrak{U}$, we have $f(x)|_{W}\notin\mathcal{F}$, and thus $f(\mathfrak{U})\subset X_{\mathcal{F}}$, which is thus nonempty.
@@
$\mathfrak{U}_{m}\cap\bigcap_{g\in W}g\cdot C^{t}_{\varphi(g)}=\varnothing\mbox{ if and only if }\mathfrak{U}\cap\bigcap_{g\in W}g\cdot C^{t}_{\varphi(g)}=\varnothing.$
Hence if $\varphi\in P_{m,t}$, we have that $\mathfrak{U}\cap\bigcap_{g\in W}g\cdot C^{t}_{\varphi(g)}\neq\varnothing$. Choosing $x\in\mathfrak{U}\cap\bigcap_{g\in W}g\cdot C^{t}_{\varphi(g)}=\bigcap_{g\in W}g\cdot U_{\varphi(g)}$ we get $f(x)|_{W}=\varphi$. Furthermore, as $f(x)\in X_{\mathcal{F}}$, we get that $\varphi=f(x)|_{W}\notin\mathcal{F}$ and thus $P_{m,t}\cap\mathcal{F}=\varnothing$.
@@ 4 Existence of universal wEDS
In this section we shall show that every finitely generated group admits a universal wEDS. Intuitively, this universal object is obtained by pasting together all possible directed sequences of effective subshifts in such a way that the resulting object is also an inverse limit. This is achieved by identifying $\mathbb{N}$ with the nodes of an unrooted universal directed tree with countably many leaves where each node carries an effective subshift and each edge carries a topological factor map between the subshifts at the nodes. In this way, every possible (infinite) sequence of subshifts and factor maps appears as an infinite path in the tree, from which we can extract a factor to any given inverse limit of effective subshifts. We remark that in this proof we do not require the indexation by the tree to be computable, the only computable objects are the subshifts on the nodes and their factor maps.
For ease of exposition, we will first do the proof in the case of a group with decidable word problem and then argue how to generalize it.
@@
Consider the collection of all tuples of the form $(h,s,X,Y,f)$ where $h$ and $s$ are non-negative integers, $X$ and $Y$ are nonempty effective subshifts on $G$ whose alphabets are finite subsets of $\mathbb{N}$ and $f$ is a topological factor map between $X$ and $Y$.
There are at most countably many effective subshifts (with their alphabets identified with finite subsets of $\mathbb{N}$) for any given group $G$. Moreover, by the Curtis-Hedlund-Lyndon theorem, there can be at most countably many topological morphisms between $X$ and $Y$ and thus at most countably many topological factor maps. Therefore the space of all such tuples is countable.
@@
Let us argue that $\mathfrak{X}$ is a wEDS. Consider the projection map $\pi_{k}\colon\mathfrak{X}\to\prod_{n\in A_{k}}X_{n}$ given by $\pi_{k}((x_{n})_{n\in\mathbb{N}})=(x_{n})_{n\in A_{k}}$ and let $\mathfrak{X}_{k}=\pi_{k}(\mathfrak{X})$. It is clear that $\mathfrak{X}_{k}$ is a subshift as it is a closed $G$-invariant subset of $\prod_{n\in A_{k}}X_{k}$. Furthermore, if we define for $i\in\mathbb{N}$ the natural projection map $g_{i}\colon\mathfrak{X}_{i+1}\to\mathfrak{X}_{i}$, then $\mathfrak{X}$ is the inverse limit of the sequence $(\mathfrak{X}_{k})_{k\in\mathbb{N}}$ endowed with the collection of factor maps $g_{i}$. Therefore, in order to show that $\mathfrak{X}$ is a wEDS, we only need to argue that for every $k\geq 0$, the subshift $\mathfrak{X}_{k}$ is effective. Indeed, let $k\geq 0$ and note first that $\prod_{n\in A_{k}}X_{k}$ is a nonempty product of effective subshifts and thus a nonempty effective subshift itself. Furthermore, by our definition of $A_{k}$, it follows that $\mathfrak{X}_{k}$ is precisely the set of elements of $\prod_{n\in A_{k}}X_{k}$ which satisfy conditions $\mathtt{C}_{n}$ for each $n\in A_{k}$. Noting that each of these conditions can be implemented with finitely many forbidden patterns (as a consequence of the Curtis-Hedlund-Lyndon theorem), we conclude that $\mathfrak{X}_{k}$ is effective.
Finally, let us show that $G\curvearrowright\mathfrak{X}$ is universal. Let $G\curvearrowright X$ be a wEDS, then it is topologically conjugate to the inverse limit of a sequence of effective subshifts. Let $(Y_{i})_{i\in\mathbb{N}}$ be said sequence of effective subshifts and for $i\in\mathbb{N}$ let $t_{i}\colon Y_{i+1}\to Y_{i}$ be the associated factor maps.
@@
    If $i<H(n)$, we choose $x_{D_{i}(n)}\in X_{D_{i}(n)}$ such that $f_{D_{i}(n)}(x_{D_{i}(n)})=x_{D_{i+1}(n)}$ and repeat the process with i1i-1 until all the values $x_{D_{i}(n)}$ are defined.
@@
Let us now consider the situation of a group $G$ which does not have decidable word problem. In this general setting, we may no longer characterize effective subshifts through forbidden patterns on $G$, but we may identify them in such a way lifting the action of $G$ to a free group [BarCarRoj2025, Proposition 7.1]. More precisely, if we let $S\subset G$ be a finite set of generators and consider the free group $F(S)$ generated by $S$ and the canonical epimorphism $\phi\colon F(S)\to G$, then a subshift $X\subset A^{G}$ is effective if and only if its pullback $\widehat{X}\subset A^{F(S)}$ is effective, where
@@
Thus in the proof of Theorem 4.1 we may consider instead of $G$ the group $F(S)$ and instead of effective subshifts $X_{n}$ and $Y_{n}$ on the group $G$, we replace them by their pullbacks $\widehat{X}_{n}$ and $\widehat{Y}_{n}$ which are subshifts on $F(S)$. The rest of the proof is exactly the same and at the end we note that the kernel of the action $F(S)\curvearrowright\mathfrak{X}$ contains $\operatorname{ker}(\phi)$ and thus the resulting induced action of $G$ is the universal $G$-wEDS. We thus obtain C.